OpenStudy (wyattp17):

Tangent is the reciprocal of which other ratio? Answer Sine Cosine Cotangent Cosecant

OpenStudy (sidsiddhartha):

Sine <--> Cosecant Cosine <--> Secant Tangent <--> Cotangent

OpenStudy (sidsiddhartha):

\[\frac{ 1 }{ \tan \theta }=\cot \theta\] \[1/\sin \theta=cosec \theta\] \[1/\cos \theta=\sec \theta\] ok got it?